OpenStudy (quinton1012):

What is 22% expressed as a fraction in simplest form?

OpenStudy (texaschic101):

22 % = 0.22 = 22/100 = 11/50

OpenStudy (yumyum247):

Yes

OpenStudy (yumyum247):

Yes

OpenStudy (quinton1012):

so 11/50 is the answer?

OpenStudy (yumyum247):

Yes

OpenStudy (yumyum247):

Yes

OpenStudy (quinton1012):

okay thanks